OpenStudy (anonymous):

can anyone help me change this to polar coordinates? integral from (0 to 3) integral from (0 to y) of xdxdy

OpenStudy (anonymous):

i have found all of the limits except for the inside upper one. i currently have...

OpenStudy (anonymous):

\[\int\limits_{\pi/4}^{\pi/2}\int\limits_{0}^{?}r^2\cos(\theta)drd(\theta)\]

OpenStudy (anonymous):

ya i have read through that